The width of a rectangle is 2 ft less then the length. The area is 3 ft ^2. Find the length and the width
Answer by mananth(16946)

let the length be x
width will be x-2
area of rectangle = L *W
x(x-2)= 3
x^2-2x=3
x^2-2x-3=0
x^2-3x+x-3=0
x(x-3)+1(x-3)=0
(x+1)(x-3)=0
x=3 OR x=-1
length cant be negative
so length = 3
width = x-2
=3-2
=1